OpenStudy (anonymous):

Find the equation of the tangent to the graph at the indicated point. f(x) = 3x + 1; a = 2 y =

myininaya (myininaya):

Have you found y'?

myininaya (myininaya):

To find the slope of our tangent line we need to evaluate the derivative at the point we wish to find the tangent line to. So by evaluating f'(2) we can find the slope of f at x=2 We can find the slope of the tangent line at x=2. So point-slope form is y-y1=m(x-x1) m is slope and (x1,y1) is a point on the line or we can say y-f(2)=f'(2)(x-2) Or y=f'(2)(x-2)+f(2)